# Bridging Online Behavior and Clinical Insight: A Longitudinal LLM-based Study of Suicidality on YouTube Reveals Novel Digital Markers

### v1

@@ 3.3 The top-down domain-expert analysis
Table 4: Psychological Factors Analysis conducted by a clinical psychologist through narrative analysis of attempt stories. The assessment comprised six sections: Risk Groups, Mental Illness, Video Motivation, Resilience Factors, Risk Factors, and Suicide Attempt Validation. The table displays the percentage of individuals in each group, with each condition determined by the psychologist based on the content described in the attempt stories. The data indicates "similar confounding" factors across most features, except for the motivations for posting the video, marked with *, and validated as significant. The table includes only factors that could be extracted from a sufficient number of users (>5%). See §A.3 for the assessment protocol and §C.2, which describes an LLM-based topic model trained on the narrative stories.
3.4 The hybrid bottom-up and top-down analysis
The bottom-up approach automatically generated LLM-based topics, which were analyzed computationally and longitudinally to identify behavioral indicators of suicide attempts (§3.1). In contrast, the top-down approach involved manual assessment of pre-defined psychological factors and analysis of suicide attempt narrative videos (§3.3). To integrate both approaches, we asked a clinical psychologist to evaluate all 166 LLM-derived topics (2.2) and flag those associated with suicidality. We then conducted the same statistical effect analysis used in the bottom-up approach on this expert-curated subset (§3.1.2).
3.4.1 What topics are known risk factors for suicide?
The clinical psychologist identified 19 topics as related to suicidality based on topic descriptions, three of which were already identified by the computational approach (§3.1). In this section, we focus on the remaining 16 topics. The full list is available in §D.
3.4.2 Temporal Effect Analysis of Suicide-Related Topics
@@
Figure 11: Anxiety and Panic Attacks. The left figure presents the average topic values across our four groups, comparing Pre- and Post-Event (suicide attempt, major life event, or synthetic condition). The right figure presents the topic’s word cloud.
[…]
As shown in Figure 11, the Attempted (During) and Control (Major Life Event) groups had lower topic prevalence before the event, with some increase observed post-event. However, post-hoc tests (Tables C.5C.6) indicated that the Attempted (Before) had a significantly higher group prevalence, and the observed temporal increase in the Attempted (During) group was not statistically significant.
Interestingly, well-established clinical suicide-related factors, such as Religion, Addiction Recovery, and Bullying (Table LABEL:tab:topic_info), did not yield statistically significant effects in our analysis, despite being flagged as relevant topics by a clinical expert. Additionally, other risk factors, such as Hopelessness or Social Isolation, did not appear in the LLM-derived topics.
4.   Discussion
This study introduces a unique, publicly available longitudinal dataset of 181 YouTube channels from individuals with verified suicide attempts, along with two additional control groups (n = 138), analyzed through a multi-method framework combining bottom-up, top-down, and hybrid approaches. Anchoring the analysis around the suicide attempt, the bottom-up computational approach identified Mental Health Struggles and YouTube Engagement as key behavioral markers, both significantly lower before the attempt compared to control groups and increasing afterward. The top-down psychological assessment of suicide narratives confirmed the suicide attempts as life-threatening, and revealed strong alignment between the narratives of Attempted (Before) and Attempted (During) groups in psychological factors such as depression and loneliness. The only notable difference was motivation to share their suicide attempt, with the latter framing it as part of a Personal Recovery Process. Lastly, the hybrid approach identified 19 out of 166 LLM-derived topics as suicide-related. However, only Anxiety & Panic Attacks was additionally supported by statistical analysis. Together, these complementary perspectives provide a richer, more nuanced understanding of how suicidality manifests on YouTube.
A particularly intriguing finding was the increase in the discussion of Mental Health Struggles after a suicide attempt. This finding is in line with previous studies indicating that mental health struggles may increase after a suicide attempt, since mental health disorders (depression, anxiety, PTSD) are often persistent or worsen post-attempts [29]. Moreover, the period after an attempt is a high-risk window for another attempt [29]. Individuals after an attempt may feel increased feelings of shame and guilt, and their interpersonal relationships, may be challenged [45]. However, this observation might appear counterintuitive based on some traditional psychological expectations, which suggest that discussions about mental health struggles typically increase prior to an attempt, as individuals at risk are more likely to seek help from mental health services [26], and may diminish post-attempt due to a sense of relief or emotional avoidance [6, 20].
Assuming the latter theory holds, possible explanations for the discrepancy between this theory and our findings may relate to the nature of the data we rely on or may be explained by psychological theory. First, consider the characteristics of the study population and platform. As noted by [26], “…alternate approaches to suicide-prevention efforts may be needed for those less likely to be seen in primary care or mental health specialty care, specifically young men.” Our research population differs from those typically studied, focusing on young people and YouTube creators in the context of their online behavior. Moreover, YouTube content may not fully capture the breadth of an individual’s emotional experiences, as video creation depends on both the technical and emotional readiness of the creator. Second, the clinical literature suggests that the increased descussion on mental health could be related to the process of post-traumatic growth [15]. Individuals who survive a suicide attempt may experience psychological growth through sharing their struggles [23]. This hypothesis is further supported by our psychological analysis, which found that the Attempted (During) group was significantly more motivated to share their stories as part of a Personal Recovery Process.
The second main finding, regarding YouTube Engagement (content aimed at driving audience interaction), is an intriguing behavioral marker, as it may reflect normal platform usage patterns rather than being directly indicative of distress. Given that creators often seek interaction to maintain their audience [49, 10, 3], this finding underscores the importance of contextualizing social media behaviors within platform norms. The finding regarding higher YouTube engagement after a suicide attempt may be explained in several ways. Engaging others may be a form of seeking connection. Individuals often feel isolated after a suicide attempt and turn to others in order to understand and validate their struggles and feelings. A social media community can provide empathy and reassurance. Another option is that the engagement is actually a cry for help [27]. This may be a way to signal that they are still struggling and need support. They might not feel comfortable asking for help directly, but hope that someone will notice and intervene.
Additional findings emerged regarding three topics identified as potential pre-attempt indicators in the computational approach: (1) LGBTQ+ Identity and Acceptance, (2) Rape and Sexual Assault, and (3) Fashion and Style. The first two align with existing literature on suicide risk factors and were independently identified as suicide-related by a clinical psychologist in the mixed-method approach. LGBTQ+ individuals are at higher risk for suicide attempts compared to the general population [43, 8]. The chronic stress experienced by LGBTQ+ individuals due to social rejection and discrimination is extremely high [46]. In addition, LGBTQ+ individuals face many barriers to mental health care [32]. Similarly, Rape and Sexual Assault are profoundly traumatic experiences that elevate the risk for attempts [12]. These traumas often lead to mental health disorders, including depression and substance abuse, which are themselves risk factors. Moreover, the victims often suffer from social isolation and stigma, which intensifies feelings of loneliness and despair. In contrast to the first two topics, Fashion and Style, a less familiar factor in suicide research, showed a significantly lower prevalence in the attempted groups. This may reflect reduced engagement with self-expression or social norms. Prior studies link between mood and clothing and suggest that fashion may indicate on positivity [33, 28]. This pattern may also align with Joiner’s Interpersonal Theory, where social withdrawal and loss of belonging contribute to suicidal behavior [19]. The decrease in fashion-related content may indicate broader patterns of social disengagement and loss of interest in activities that typically foster interpersonal connection. However, given that Fashion and Style are not conventionally examined in suicide research, further studies are needed to clarify this relationship.
Despite these promising findings, this study has several limitations. First, the sample is inherently biased toward individuals who are active on YouTube, which may not be representative of the broader population at risk for suicide. Additionally, the reliance on public content means that more private or intimate indications of distress may not be captured (as would be the case in, e.g., clinical records and personal messages). Another limitation is that all clinical annotations were performed by a single expert, which may introduce subjective bias. This limitation stems from the demanding nature of the task, which required reviewing and coding hours of suicide narratives. Future studies should consider incorporating data from other platforms and employing methods to assess more private, less curated expressions of suicidality.
In summary, our study adds to the growing body of literature on social media and mental health while also advancing the use of YouTube for suicide research. By integrating qualitative and quantitative methods and emphasizing longitudinal analysis considering suicide attempts, we highlight the significance of a temporal approach compared to broader group-based analyses. We hope this work encourages further interdisciplinary research in this critical area.
Appendix A Appendix A - Data Collection
A.1 Inclusion and exclusion criteria
Stage (Scope) Filtering Step Group
Search (Video) Search Queries (OR)
“my suicide attempt”
“my suicide story”
“I am a suicide attempt survivor”
“attempted suicide”
“take my life”
“my relative died”
“losing my relative
“I lost my relative
[…]
father, mother}
Search (Video) Category
Excluded query videos from the following categories:
["Gaming", "Sports", "Comedy", "Music", "News & Politics",
Preprocessing (Video) File Size Limit Not applied Excluded $>$50MB
Preprocessing (Transcript) Transcripts Excluded transcripts if: (1) $<$3 sentences, (2) $>$50% repeated (hallucinated) content, (3) $>$30% non-English, or (4) $>$2 speakers. Videos which weren’t filtered are considered valid.
Preprocessing (Channel) Activity Threshold Excluded channels with $<$10 valid videos
Final Sample Size n=42 n=139 n=43 n=95
As part of the data collection, we used the YouTube API ,Google YouTube API Documentation, to search for channels of people who are relevant for our research, i.e., those who attempted suicide - Attempted (During) and Attempted (Before), and those who experienced a major life event - Control (Major Life Event). See Table A.1 for full inclusion and exclusion criteria for the research groups.
A.2 Data preprocessing:
During the data collection §2.1.1, we implemented the following preprocessing steps:
  1. 1.
    Downloading: We downloaded all channel uploads exceeding 60 seconds (no upper limit on video duration), excluding unavailable videos (e.g., private uploads). The downloading process, which spanned several months, utilized the yt-dlp package to obtain videos in .mp3 format and bright-data services for port forwarding services.
  2. 2.
    Whisper Transcription: We used the Whisper (medium) [41], model for audio-to-text conversion with specific parameters to enhance transcription accuracy (no speech threshold=0.3, condition on previous text=True).
  3. 3.
    1. (a)
      Too short - fewer than three sentences.
    2. (b)
      Contain hallucinations - repeated words or sentences comprising more than 50% of the document, often resulting from non-verbal content like background music [50].
    3. (c)
      Non-English Content - contain more than 30% non-English content (detected automatically).
    4. (d)
      Non-Personal - more than 2 speakers (detected automatically).
    The remaining videos were labeled as valid.
  4. 4.
    Transcript Segmentation: To accommodate the token limits of modern embedding models like Sentence-BERT [42] (which has a 512-token threshold), we divided transcripts into chunks of a maximum of 80 tokens, ensuring each chunk contained complete sentences. This segmentation allows for better representation of multiple topics within a single video.
A.3 Human evaluation
As part of our annotation process, we used the Label Studio [47] platform to create two protocols. In both, the annotators were provided with a YouTuber image, Link to the YouTuber’s channel, Attempt/Major Life Event narrative’s video, accompanied with audio, and transcription of the video, and summary (automatically generated by GPT-4 [36]).
The annotators were asked to answer the following questions:
  • Relevance -
  • Demographic Information - age, ethnicity, gender.
  • Event information -
    • Does the story talk about a suicide attempt/major life event?
    • Who did the attempt/died?
    • What is the channel owner’s age at the time of the event?
  • Risk Factors Information:
    • What risk factors for suicide are present (e.g., depression)?
    • What resilience factors are evident (e.g., family support)?
  • Clinical Information
Figure A.1: The annotation interface from Label-Studio [47], displaying a partial example provided to the annotators. The interface supported two main annotation processes: (1) Non-Expert Assessment, which involved classifying personal channels, estimating demographic information, assessing the relevance of the story, extracting event-related details (e.g., suicide attempt date); and (2) Psychological assessment of suicide attempt narratives (conducted by a clinical psychologist), which included evaluating risk factors, risk groups, mental health illness, and distinguishing suicide attempts from non-suicidal self-harm. Full example of annotation interface is attached in a separate file.
A.4 Matching procedure:
Overview
(3) The first step used propensity score-based matchingA.4.3) [7], which estimates similarity based on numerical and categorical features using a logistic regression model trained on demographic attributes (e.g., age) and channel-related features (e.g., number of uploads). Matching was performed using k-nearest neighbor on the resulting propensity scores.
(4) The second step involved LLM-based refined matchingA.4.4), leveraging GPT-4 [36]. This step incorporated channel content, aiming to control for general themes reflected in the textual content of the channel (e.g., gaming, coding, fashion). Matches were scored on a 1–5 scale based on textual content descriptions, and only channels rated 3 or higher were retained.
Technical Details
A.4.1 Queries based on treatment group
A.4.2 Automatic and manual filtering personal channels
A.4.3 Propensity Score-Based Matching:
Propensity Score Estimation:
  • Demographic features: gender and age (manually annotated).
Propensity Score Matching:
A.4.4 LLM-based refined matching:
Lastly, we incorporated textual features, channel descriptions, upload titles, and video descriptions, to further refine the matches using GPT-4 [36]. For cost and efficiency, three uploads per channel were preprocessed and summarized using GPT-4. Then, we used GPT-4 again to evaluate and rate each match on a scale of 1 to 5 (using the prompt described in B.1.2), providing explanations for its choices. Matches rated 3 or higher were retained, as they were deemed to be of sufficient quality for analysis. Matches rated below 3 were excluded to reduce noise and improve the reliability of the results.
A.5 Dateset statistics of pre- and post event
The dataset comprised four distinct groups of YouTubers, with varying sample sizes and upload patterns before and after the reference event, described fully in Table A.2. The treatment group, Attempted (During), consists of 42 creators who attempted suicide, of whom 38 had valid upload content after the attempt.
Figure A.2: Box plots representing the number of days between the last upload before the suicide attempt (left) or major life event (right), to the event’s themselves, and afterwards. The y-axis represents logarithmic scale.
In addition, we analyzed the number of days between the last upload before the major life event or suicide attempt, the event itself, and the first upload after the event, in Tables A.3 and A.4. The Attempted (During) group, as seen in Table A.3 has a median time between the last upload before the attempt and the first upload after the attempt of 20 days, while the Control (Major Life Event) group, as seen in Table A.4 had 6 days between the two. To better understand the distribution of inactivity periods, we visualized the gaps using boxplots with a log-scaled y-axis (Figure A.2). While the Attempted group displayed longer gaps overall, we observed that both groups had similarly skewed distributions with multiple long-tail outliers.
We further compared each type of interval between the Attempted (During) and Control (Major Life Event) groups using Welch’s t-test, based on the absolute duration in days and excluding missing values. The analysis revealed that the differences in all three interval types were not statistically significant, indicating similar variability in posting behavior across the two groups.
@@ A.5 Dateset statistics of pre- and post event
Mean Median Std Mean Median Std
@@
Attempted (Before) Middle of Uploads 139 102 44 159 98 39 153
midpoint of upload dates was determined
by sorting the unique dates and selecting
the median as the reference point

### v2

@@ C.3 Supplementary Figures and Tables
Table C.1: Psychological Factors Analysis conducted by a clinical psychologist through narrative analysis of attempt stories. The assessment comprised six sections: Risk Groups, Mental Illness, Video Motivation, Resilience Factors, Risk Factors, and Suicide Attempt Validation. The table displays the percentage of individuals in each group, with each condition determined by the psychologist based on the content described in the attempt stories. The data indicates "similar confounding" factors across most features, except for the motivations for posting the video, marked with *, and validated as significant. The table includes only factors that could be extracted from a sufficient number of users (>5%). See §A.3 for the assessment protocol and §C.2, which describes an embedding-based with LLM support topic model trained on the narrative stories.
@@
Figure C.5: Anxiety and Panic Attacks. The left figure presents the average topic values across our four groups, comparing Pre- and Post-Event (suicide attempt, major life event, or synthetic condition). The right figure presents the topic’s word cloud.
Figure C.6: Longitudinal Comparison of YouTube Engagement (left) and Mental Health Struggles (right) relative to $t=0$ in the Attempted (During) Group. The figures depict temporal trends of these topics within the Attempted (During) group, with the timeline aligned to the suicide attempt. At each time step, comparisons were made to $t=0$ using the paired t-test. Asterisks (*) indicate statistically significant differences. The full statistical procedure is detailed in §B.3.
Appendix D Appendix D - Statistical and Robustness Analyses
This appendix presents the full statistical analyses supporting the findings described in the results sections and Appendix §C (Supplementary Results). It includes mathematical formulation of the GLMM model (§D.1), and a series of robustness and validation analyses (§D.2) that examine potential confounding factors, engagement metrics, and upload activity patterns across groups. Together, these results quantify and validate the effects reported, ensuring that the observed behavioral and linguistic differences are statistically reliable and methodologically sound.
D.1 Effects Analysis
As part of both the bottom-up and hybrid approaches, we estimated a generalized linear mixed model (GLMM) using a Beta likelihood and a logit link function to model the bounded response $y_{ij}\in(0,1)$, corresponding to the mean topic probability for user $i$ at time $j$. We calculated the aggregated mean topic probabilities before and after the reference event. The model included fixed effects for Time (0 = Before, 1 = After), Group (Attempted (During), Attempted (Before), Control (Matches), Control (Major Life Event)), and their interaction (Group $\times$ Time), together with continuous and demographic covariates (Age, Gender, Ethnicity, Upload activity). Random intercepts were specified at the user level to account for repeated measures. The Attempted (During) group and the pre‐event period (Before) served as reference categories. Thus, each coefficient represents the change in logit‐transformed topic prevalence relative to this baseline. Group effects were parameterized as contrasts against the Attempted (During) group:
$\text{logit}\big(E[y_{ij}]\big)=\beta_{0}+\beta_{1}\,\text{Time}_{ij}+\sum_{g=1}^{G-1}\beta_{2g}\,\text{Group}_{g,i}+\sum_{g=1}^{G-1}\beta_{3g}\,(\text{Group}_{g,i}\times\text{Time}_{ij})+\boldsymbol{\beta_{4}}^{\top}\mathbf{X}_{ij}+b_{0i},$ (1)
where $\mathbf{X}_{ij}$ represents the matrix of user‐level covariates, and $b_{0i}$ denotes user‐specific random intercepts, assumed to follow $b_{0i}\sim\mathcal{N}(0,\sigma_{b}^{2})$.
D.1.1 Temporal Effects: Mental Health Struggles and YouTube Engagement
For the Mental Health Struggles topic, the Attempted (During) group demonstrated a significant increase in topic prevalence following the event ($\beta=0.55$, $z=3.82$, $p_{adj}<.001$), corresponding to a 74% increase in the odds of mental health content post-event. The Attempted (Before) group showed a significantly smaller temporal increase relative to the Attempted (During) group ($\beta=-0.44$, z=2.72z=-2.72, $p_{adj}=.016$), indicating that their post-event change was attenuated compared to the reference group. The control groups also showed patterns of smaller temporal increases (Control (Matches): $\beta=-0.36$, z=2.08z=-2.08, $p=.038$; Control (Major Life Event): $\beta=-0.40$, z=1.99z=-1.99, $p=.046$)
The YouTube Engagement topic exhibited a significant post-event increase in the Attempted (During) group ($\beta=0.52$, $z=4.80$, $p_{adj}<.001$), reflecting a 67% increase in the odds of YouTube engagement content. All comparison groups demonstrated significantly smaller temporal increases relative to the Attempted (During) group. The Attempted (Before) group showed a significantly attenuated post-event increase ($\beta=-0.35$, z=2.90z=-2.90, $p_{adj}=.016$). The Control (Matches) group demonstrated the most pronounced differential effect ($\beta=-0.53$, z=4.13z=-4.13, $p_{adj}<.001$), with substantially less temporal change than the reference group. The Control (Major Life Event) group similarly exhibited a significantly smaller post-event increase ($\beta=-0.42$, z=2.78z=-2.78, $p_{adj}=.027$) compared to the Attempted (During) group.
D.1.2 Baseline Group Differences: YouTube Engagement and Fashion and Style
Independent of temporal changes, baseline differences in YouTube Engagement were observed across groups. The Control (Matches) group ($\beta=0.83$, $z=6.05$, $p_{adj}<.001$) and Control (Major Life Event) group ($\beta=0.76$, $z=3.75$, $p_{adj}=.001$) both exhibited markedly higher baseline prevalence compared to the Attempted (During) group, corresponding to 129% and 115% higher odds, respectively. Male users were associated with significantly lower baseline engagement ($\beta=-0.31$, z=2.84z=-2.84, $p_{adj}=.023$), corresponding to 27% lower odds relative to female users.
The Fashion and Style topic showed significant group and gender differences at baseline, with no significant temporal changes detected. The Control (Matches) group demonstrated higher prevalence of fashion-related content compared to the Attempted (During) group ($\beta=0.89$, $z=2.86$, $p_{adj}=.01$), representing 144% higher odds. Male users showed significantly lower engagement with this content ($\beta=-0.66$, z=2.55z=-2.55, $p_{adj}=.026$), with 48% lower odds compared to female users.
D.2 Robustness and Validation Analyses
To evaluate whether the observed pre-attempt dip and post-attempt rise in engagement could be explained by non-behavioral factors, we conducted a series of robustness checks addressing potential factors. Specifically, we examined (1) sensitivity analysis to ambiguous cases; (2) platform-level shifts that might reflect external events such as the COVID-19 pandemic; (3) group-level differences in audience engagement or channel characteristics; and (4) upload-frequency biases at the individual channel level.
D.2.1 Sensitivity Analysis
To assess the robustness of our findings to ambiguous cases, we performed a sensitivity analysis excluding 14 users (out of 139 Attempted (Before); $13.8\%$) who posted content describing a suicide attempt but were clinically evaluated as having only suicidal thoughts (see Table C.1). Notably, this exclusion removed 6 male users from the sample. After excluding these users, all main effects and temporal patterns remained consistent with the primary analysis (Table 3; GLMM results), except for Group difference to Attempted (Before), which were no longer significance in any topic ($p_{adj}>0.05$). These results indicate that our primary conclusions regarding topic prevalence and group × time interactions are robust to this source of ambiguity, while the Group difference to specific groups should be interpreted with caution due to its sensitivity to sample composition.
D.2.2 Control Measures Analysis (Platform-Level Effects)
To ensure that the temporal patterns identified around the suicide attempt were not artifacts of global platform dynamics, we examined topic trends using an external, time-anchored reference event unrelated to individual users—the onset of the COVID-19 pandemic in the United States (March 1, 2020). The same temporal alignment and analysis procedures as in B.2 were applied to all four research groups.
Figure D.1: Temporal dynamics of topic prevalence around two reference points: the onset of the COVID-19 pandemic in the United States (March 1, 2020; left) and the individual-level reference events (suicide attempt, major life event, or midpoint of upload; right). Each panel displays three topics—COVID-19 Pandemic (top) and Mental Health Struggles (middle), YouTube Engagement —illustrating that the global COVID-19 event elicited parallel changes across all groups, whereas the attempt-related dynamics exhibit distinct within-group patterns.
As shown in Figure D.1, both the treatment and control groups exhibited an expected increase in the topic COVID-19 Pandemic around March 2020, consistent with a global shift in YouTube content during that period. Crucially, there were no significant between-group differences in this topic, indicating that all groups were equally affected by this exogenous event. Moreover, while a gradual increase was observed in the topic Mental Health Struggles during the pandemic, its temporal dynamics differed markedly from those observed around the suicide attempt event.
D.2.3 Engagement Metrics Analysis (Audience and Channel Composition Effects)
Next, we tested whether the engagement pattern could be explained by audience composition or content popularity rather than behavioral change. To this end, we compared aggregate engagement metrics across all four research groups, including the number of uploads, likes, comments, and average video duration. These measures were analyzed both cross-sectionally and longitudinally relative to each group’s reference event.
Figure D.2: Box plots showing the distribution of uploads, likes, comments, and video duration, aggregated across all channel videos. Each box-plot represents the distribution across users in each group.
As illustrated in Figure D.2, the distributions of engagement metrics for the Attempted (During), Attempted (Before), and Control (Matches) groups were largely similar. The Control (Major Life Event) group showed somewhat higher mean values for uploads, likes, and comments, reflecting greater general channel activity. A temporal comparison of likes over time (Figure D.3) further confirmed that these differences were stable, rather than reflecting dynamic changes around the event period.
Figure D.3: Number of likes over time around the reference event, comparing the four research groups.
D.2.4 Activity Analysis (Upload Frequency)
Finally, to assess whether the engagement pattern could be explained by disproportionate posting activity, we analyzed upload behavior at the channel level. Specifically, we examined (a) temporal gaps between uploads before and after the event, and (b) total number of uploads per user in the pre- and post-event periods. This analysis focused on the Attempted (During) and Control (Major Life Event) groups, as both included individually dated reference events (in contrast to synthetic midpoints used for the other groups).
Figure D.5 summarizes the temporal gaps between uploads for both groups. We analyzed the number of days between the last upload before the reference event (suicide attempt or major life event), the event itself, and the first upload after the event. In the Attempted (During) group, the median interval between the last and first upload surrounding the attempt was 20 days, whereas in the Control (Major Life Event) group it was 6 days. Although some numerical differences were observed, Welch’s t-tests comparing the three interval types (before-to-event, event-to-after, and before-to-after), excluding missing values, revealed no statistically significant group differences, indicating similar variability in posting behavior across the two groups.
We also compared the total number of uploads before and after the reference event (Figure D.4). Across all groups, upload frequency remained relatively stable, with the Control (Major Life Event) group showing slightly higher activity, consistent with the elevated engagement metrics observed in §D.2.3.
@@ D.2.4 Activity Analysis (Upload Frequency)
Mean Median SD Mean Median SD
@@
Attempted (Before) Midpoint of Uploads 139 102 44 159 98 39 153
Reference event corresponds to the midpoint
of the upload period, determined by sorting
unique upload dates and selecting the median
Control (Major Life Event) Major Life Event 43 192 88 330 119 50 132
Examples include death or suicide attempt
of a close relative (parent, child, sibling)
Control (Personal) Midpoint of Uploads 95 78 42 99 76 42 97
Synthetic reference midpoint calculated using
the median of upload dates
Figure D.4: Pre- and post-event upload activity across research groups. Top figure visualizes boxplots total uploads per user before and after the reference event, and the bottom summarizes corresponding descriptive statistics and methodological notes.